OpenStudy (anonymous):

Please help! Given: tan theta=-5/12 and theta is in Quadrant IV. Sketch an angle in standard position that satisfies the given conditions. The correct graph is??

OpenStudy (apoorvk):

fourth quadrant, and tan theta = -5/12 = y/x So, plot (12,-5) on graph and connect it with the origin. The angle formed with the positive X-axis would be your required angle.

OpenStudy (anonymous):

Thanks!